The coordinates of the turning point of the graph y-2x^2=-4x+1 are
Im pretty sure the answer is (1,1) but i don't know for sure and plus can you show how you got the answer, thank you!

Answer by lwsshak3(11628)

y-2x^2=-4x+1
Finding coordinates of the turning point in a parabola is the same as finding the coordinates of the vertex.
There are two methods you can use.
First, change the equation to this form,
y=2x^2-4x+1
a=2,b=-4,c=1
the x-coordinate is equal to -b/2a = -(-4)/2*2=1
the y-coordinate is then equal to 2*1^2-4*1+1=-1
the other method is to complete the square into the form, y=A(x-h)^2+k, h,k being the coordinates of the vertex
2(x^2-2x+1)+1-2=2(x-1)^2-1
from this form you can read that the coordinates of the vertex are (1,-1)
The answer you got was almost correct. You just missed a sign which we all do now and then.